OpenStudy (anonymous):

solve the equation on the interval 0

OpenStudy (anonymous):

you know a number whose cosine is \(\frac{\sqrt{3}}{2}\) ?

OpenStudy (anonymous):

should look kind of familiar

OpenStudy (anonymous):

\[\pi/6 , 11\pi/6\]

OpenStudy (anonymous):

ok so if \(2\theta=\frac{\pi}{6}\) then that makes \(\theta =\frac{\pi}{12}\) that is one of the answers, there are more

OpenStudy (anonymous):

so then \[11\pi/12\] would be the second?

OpenStudy (anonymous):

that would be the second one, but there may be more

OpenStudy (anonymous):

because if you go around the unit circle again, you get to \[\frac{13\pi}{6}\] which is not in the interval \([0.2\pi]\) but half of it is

OpenStudy (anonymous):

would tou also do cos= \[-\sqrt{3}/2\]

OpenStudy (anonymous):

no, it doesn't say anything about that

OpenStudy (anonymous):

but if \[2\theta =\frac{13\pi}{6}\] then \[\theta=\frac{13\pi}{12}\]which is in your interval

OpenStudy (anonymous):

then 23pi/12 can also be an option cause its in that interval